OpenStudy (anonymous):

3/48 is what part of 1/2

OpenStudy (paki):

3/48= x*(1/2) x=(3/48)*2 solve foe "X"

OpenStudy (anonymous):

sorry it is 3/48 is what part of 1/12

OpenStudy (paki):

3/48= x*(1/12) x=(3/48)*12... now solve foe "X"

OpenStudy (anonymous):

ok thnx